OpenStudy (anonymous):

sin(theat/2)=1/2 How would you solve this???

OpenStudy (turingtest):

you have to know your special values for trig functions what is arcsin(1/2) ?

OpenStudy (anonymous):

That would be 60. Because the question is asking sin of what divided by a half equals 30. Since sin30 = 1/2; 30 x 2 = 60; And therefore theta is 60 because sin (60/2) is equal to 1/2.

OpenStudy (turingtest):

first use the inverse trig function arcsin:\[\sin(\frac\theta2)=\frac12\implies\frac\theta2=\sin^{-1}(\frac12)\]so at what angle x is sinx=1/2 ?

OpenStudy (anonymous):

pi/6=x what is a general formula?

OpenStudy (turingtest):

to find x in the above question? there isn't one, you just got lucky that 1/2 is a special value of sinx so is sinx=sqrt3/2 and such if it were sinx=0.27368... or something funky like that, you would have ot break out the calculator and use the arcsin function